### v2

# AHA! Animating Human Avatars in Diverse Scenes with Gaussian Splatting

Abstract
We present a novel framework for animating humans in 3D scenes using 3D Gaussian Splatting (3DGS), a neural scene representation that has recently achieved state-of-the-art photorealistic results for novel-view synthesis but remains under-explored for human-scene animation and interaction. Unlike existing animation pipelines that use meshes or point clouds as the underlying 3D representation, our approach introduces the use of 3DGS as the 3D representation to the problem of animating humans in scenes. By representing humans and scenes as Gaussians, our approach allows for geometry-consistent free-viewpoint rendering of humans interacting with 3D scenes. Our key insight is that the rendering can be decoupled from the motion synthesis and each sub-problem can be addressed independently, without the need for paired human-scene data. Central to our method is a Gaussian-aligned motion module that synthesizes motion without explicit scene geometry, using opacity-based cues and projected Gaussian structures to guide human placement and pose alignment. To ensure natural interactions, we further propose a human–scene Gaussian refinement optimization that enforces realistic contact and navigation. We evaluate our approach on scenes from Scannet++ and the SuperSplat library, and on avatars reconstructed from sparse and dense multi-view human capture. Finally, we demonstrate that our framework allows for novel applications such as geometry-consistent free-viewpoint rendering of edited monocular RGB videos with new animated humans, showcasing the unique advantage of 3DGS for monocular video-based human animation. To gauge the full quality of our results we urge the reader to watch the supplementary available at https://miraymen.github.io/aha/
footnotetext: *Work done as intern at Snap. \$Work done while at Snap
1 Introduction
Human animation in 3D scenes is essential for applications ranging from video gaming and computer-generated imagery (CGI) to robotics. Recent research has made significant progress on generating humans in 3D scenes (hassan2019prox; jiang2024scaling; hassan_samp_2021; zhao2023dimos; hwang2025scenemimotioninbetweeningmodeling), where humans are typically represented as either 3D skeletons or meshes, while background scenes are represented as meshes or point clouds. These representations are compact and versatile, capable of modeling a wide variety of surfaces. However, a fundamental limitation persists: there is almost always a domain gap between rendered results and real images due to limitations in lighting, materials, and geometric fidelity.
In parallel, neural scene representations have emerged, beginning with NeRF (mildenhall2020nerf) and recently 3D Gaussian Splatting (kerbl3Dgaussians), enabling photorealistic rendering of objects, humans, and full 3D scenes from novel viewpoints or in novel poses. Yet, despite their success in rendering quality, neural representations have seen little to no adoption in human-scene animation pipelines, which continue to rely on mesh and point cloud–based frameworks.
Gaussian Splatting as a 3D representation for human-scene animation, in theory offers natural advantages over existing mesh based representations: First, 3DGS enables photorealistic rendering of human-scene interactions with superior lighting and material fidelity. Second, 3DGS allows for reconstructing scenes with only a monocular video (ling2024dl3dv) captured from a mobile phone, thus allowing for applications such as geometry consistent free viewpoint rendering of videos with new humans, personalized content creation and gaming in scenes from mobile-captured videos. Such applications are difficult with a mesh based representation as estimating meshes, or pointclouds from only monocular scene videos remains challenging (wang2025vggt). This motivates the central question addressed in this paper: Can neural scene representations—specifically Gaussian Splatting—be effectively used as a 3D representation for human animation in 3D scenes? (Fig. LABEL:fig:teaser).
Several obstacles prevent a direct extension of 3DGS to human animation in 3D scenes. First, most existing work on human–scene interaction synthesis (hassan_samp_2021; jiang2024scaling; hwang2025scenemimotioninbetweeningmodeling; zhao2023dimos) assumes paired human motion data with scene geometry. Such datasets are difficult to collect at scale, and conversion from mesh-based annotations into Gaussians is non-trivial. Second, human-scene animation requires motion synthesis that respects both the structure of the scene and the natural dynamics of the human pose, which for a non-mesh representation remains non-trivial. Furthermore, unlike meshes, 3DGS does not provide explicit topology or clean geometry, complicating tasks like surface-based contact modeling.
To address these challenges, we offer a novel perspective for human–scene animation, grounded in two key insights: First, rendering of humans and scenes in 3DGS can be decoupled from motion synthesis. That is, we can reconstruct humans and scenes independently, animate humans in a canonical space, and then place them back into reconstructed 3DGS scenes. This is common in classical graphics pipelines for meshes, where canonical models are animated via skinning or rigging, and has recently been adopted for animatable 3DGS avatars as well. However, prior work has primarily studied such Gaussian avatars in isolation. In contrast, our contribution is to extend this paradigm to human–scene animation, where avatars must not only be animated but also consistently placed and rendered inside reconstructed 3DGS scenes. Second, motion synthesis can itself be decoupled from explicit geometry: even though 3DGS does not provide watertight surfaces, we show that its opacity fields and projected Gaussian structures offer sufficient cues to guide human placement.
Our framework proceeds in two stages. First, we reconstruct humans as animatable Gaussians from either multiview capture using an off-the-shelf module. These Gaussians are then posed using a Gaussian-aligned motion module (Sec. 3.2), which computes scene-aligned motion parameters by relying on opacity-based culling and orthographic projection of scene Gaussians for path finding. Our core contribution is the migration and adaptation of traditional scene-mesh human interaction techniques (including RL-based locomotion and motion transitions) to 3DGS. To further ensure realistic interactions, we introduce a human–scene Gaussian refinement optimization (Sec. 3.3) that adjusts the placement and motion of humans for natural contact and navigation within the scene.
To showcase the applicability of our method on diverse datasets, we present results on scenes from the Scannet++ dataset (yeshwanthliu2023scannetpp) and from scenes downloaded from the publicly available SuperSplat 3DGS library (supersplat). To demonstrate the efficacy of our method on different Avatar reconstruction datasets, we also showcase results on Avatars from the BEHAVE (bhatnagar22behave) and Avatarrex datasets (zheng2023avatarrex). The results from BEHAVE demonstrate that our method works on avatars reconstructed from sparse camera setups. We finally demonstrate the utility of our presented framework for geometry consistent free viewpoint rendering of monocular videos with new animated humans on several monocular videos from the DL3DV dataset (ling2024dl3dv), showcasing the unique advantage of 3DGS for casual video-based human animation.
To summarize our contributions are as follows:
  • We introduce the 3D Gaussian Splatting representation to the classical Computer Graphics problem of animating humans in 3D environments
  • We demonstrate that our framework can be used for geometry consistent free viewpoint rendering of monocular videos edited with new animated humans
  • We introduce a novel Gaussian aligned motion module for motion synthesis in scenes represented as 3D Gaussians
  • We introduce a human scene Gaussian refinement optimization for correct placement of human Gaussians in scenes represented using 3DGS leading to better contact and interactions.
2 Related Work
Neural Rendering Following the publication of NeRF (mildenhall2020nerf), there has been significant research on Neural Rendering  (nerf_review). Nerf is limited by its computational complexity and despite several follow-up improvements (mueller2022instant; barron2022mip; barron2023zipnerf; nerfstudio), the high computational cost of NeRF rermain. 3DGS introduced in (kerbl3Dgaussians) addresses this limitation by representing scenes with an explicit set of primitives shaped as 3D Gaussians, extending previous work  (lassner2021pulsar). 3DGS rasterizes Gaussian primitives into images using a splatting algorithm (westover1991phdsplatting). 3DGS originally designed for static scenes has been extended to dynamic scenes  (shaw2023swags; luiten2023dynamic; Wu2024CVPR; lee2024ex4dgs; li2023spacetime), slam-based reconstruction, (keetha2024splatam), mesh reconstruction (Huang2DGS2024; guedon2023sugar) and NVS from sparse cameras (mihajlovic2024SplatFields).
Human Reconstruction and Neural Rendering Mesh-based templates (SMPL-X:2019; smpl2015loper) have been used to recover 3D human shape and pose from images and video (Bogo2016keepitsmpl; kanazawaHMR18). However, this does not allow for photoreal renderings. In (alldieck2018video; alldieck19cvpr) recover a re-posable human avatar from monocular RGB. However their use of a mesh template also does not allow for photorealistic renderings. Implicit functions (mescheder2019occupancy; park2019deepsdf) have also been utilized to reconstruct detailed 3D clothed humans (chen2021snarf; alldieck2021imghum; saito2020pifuhd; he2021arch++; huang2020arch; deng2020nasa). However, they are also unable to generate photorealistic renderings and are often not reposable. Several works (peng2021animatable; guo2023vid2avatar; weng_humannerf_2022_cvpr; jian2022neuman; haberman2023hdhumans; heminggaberman2024trihuman; li2022tava; liu2021neural; xu2021hnerf) build a controllable NeRF that produces photorealistic images of humans from input videos. Unlike us, they do not model human-scene interactions. With the advent of 3DGS, several recent papers use a 3DGS formulation (kocabas2023hugs; qian20233dgsavatar; moreau2024human; abdal2023gaussian; zielonka2023drivable; moon2024exavatar; li2024animatablegaussians; pang2024ash; lei2023gart; hu2024gaussianavatar; li2024animatable2; zheng2024gpsgaussian; jiang2024robust; dhamo2023headgas; qian20233dgs; xu2024gaussian; junkawitsch2025eva) to build controllable human or face avatars. Unlike our method, they do not model human-scene interactions. Prior works have also extended the 3DGS formulation to model humans along with their environment, (xue2024hsr; tomie; mir2025gaspacho), but unlike us, they either do not focus on human animation in 3D scenes.
Humans and Scenes Human-scene interaction is a recurrent topic of study in computer vision and graphics. Early works (fouhey2014people; wang2017binge; gupta20113d) model affordances and human-object interactions using monocular RGB. The collection of several recent human-scene interaction datasets (Hassan2021-gb; mir20hps; hassan2019prox; savva2016pigraphs; taheri2020grab; bhatnagar22behave; jiang2024scaling; cheng2023dnarendering; zhang2022couch) has allowed the computer vision community to make significant progress in joint 3D reconstruction of human-object interactions (xie2022chore; xie2023vistracker; xie2024template; zhang2020phosa). These datasets have also led to the development of methods that synthesize object conditioned controllable human motion (zhang2022couch; starke2019neuralstate; hassan21cvpr; diller2023cghoi). All these methods represent humans and scenes as 3D meshes and inherit the limitations of mesh-based representations including their inability to generate photorealistic images, while our method allows for photorealistic renderings of humans and scenes.
Human Animation Human animation is another extensively studied problem in vision and graphics. Motion matching (reitsma2007evaluating), learned motion matching (clavet16motionmatching; holden2020learned) and motion graphs  (lee2002interactive; fang2003efficient; kovar2008motion; safonova2004synthesizing; Safonova:2007:InterpolatedGraphs) are common methods employed in the video-gaming industry for generating kinematic motion sequences. Deep learning variants (holden2017phase; starke19neural; starke21martialarts; starke20local) have also gained popularity. Diffusion Models (tevet2023human) have emerged as a powerful paradigm for human motion synthesis. Several follow-up works extend the original Motion Diffusion model with physics(yuan2023physdiff), blended-positional encoding (barquero2024seamless), and for fine-grained controllable motion synthesis (karunratanakul2023dno; pinyoanuntapong2024controlmm; xie2024omnicontrol). Reinforcement learning (zhang2022wanderings; zhao2023dimos) is another oft-used paradigm used for motion synthesis. Diffusion models have also been used as latent-motion models (Zhao:DartControl:2025) but unlike us, they only focus on clean, noise-free mesh-based scene representations and ignore the rendering aspects of human-scene interaction leading to limited rendering quality and photorealism.
Figure 1: Using a latent motion model and 3DGS scene representation, we synthesize human motion that confirms with the 3D scene Gaussians using Gaussian-aligned motion module. We adapt RL based controllers and latent optimization for 3DGS scenes. We further refine these Gaussians for correct placements and contact. These composited human and scene Gausssians can be rendered from any viewpoint to generate photoeal images.
3 Method
We present a method that enables virtual humans represented using Gaussian splats to navigate and interact in complex environments reconstructed as 3D Gaussian scenes. Our framework consists of three key components: (1) Gaussian Reconstruction: We reconstruct both scenes and humans as 3D Gaussians from RGB images. For scenes, we use standard 3DGS reconstruction, while for humans, we learn human Gaussian representations that can be animated with different SMPL poses (Sec. 3.1). (2) Gaussian-Aligned Motion Synthesis: Central to our approach is a novel Gaussian-aligned motion module (Sec. 3.2), which uses the reconstructed scenes (Sec. 3.1) and a latent-variable based motion synthesis framework (using RL and latent space optimization adapted to 3DGS) to synthesize motion parameters aligned with the 3DGS scenes (3) Differentiable Contact Refinement in 3DGS: We use the synthesized human motion data to animate human Gaussians and apply a novel refinement algorithm for correct human-scene interaction (Sec. 3.3) (Fig. 2). The refinement module detects contact frames from motion data and optimizes translation vectors to enforce proper contact between human and scene Gaussians while maintaining temporal smoothness and avoiding penetrations. These composed Gaussians can be rendered from any camera viewpoint to produce videos of humans interacting with diverse scenes. Figure 1 provides an overview.
3.1 Gaussian Reconstruction
Scene Gaussians. Given monocular or multi-view video of a static scene, we model the environment as a set of $N_{S}$ anisotropic 3D Gaussians $\mathcal{G}^{S}=\{(\bm{\mu}_{i},\bm{\Sigma}_{i},\alpha_{i},\mathbf{c}_{i})\}_{i=1}^{N_{S}}$, where each Gaussian has center $\bm{\mu}_{i}\in\mathbb{R}^{3}$, covariance $\bm{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$, opacity $\alpha_{i}\in(0,1)$, and possibly view-dependent color $\mathbf{c}_{i}$. Under the camera projection $\Pi_{t}$, each Gaussian projects to an ellipse with screen-space covariance $\bm{\Sigma}^{2D}_{i,t}=J_{t}\bm{\Sigma}_{i}J_{t}^{\top}$, where $J_{t}=\left.\frac{\partial\Pi_{t}}{\partial\mathbf{x}}\right|_{\bm{\mu}_{i}}$ and $\mathbf{u}_{i}=\Pi_{t}(\bm{\mu}_{i})$. Its pixel contribution at $\mathbf{u}$ is $g_{i,t}(\mathbf{u})=\exp\!\big(-\tfrac{1}{2}(\mathbf{u}-\mathbf{u}_{i})^{\top}(\bm{\Sigma}^{2D}_{i,t})^{-1}(\mathbf{u}-\mathbf{u}_{i})\big)$, yielding effective opacity $\hat{\alpha}_{i,t}(\mathbf{u})=\alpha_{i}\,g_{i,t}(\mathbf{u})$. The rendered image is obtained via front-to-back alpha compositing,
$\hat{\mathbf{I}}_{t}(\mathbf{u})=\sum_{i\in\mathcal{S}_{t}(\mathbf{u})}\Big(\prod_{j\in\mathcal{S}_{t}(\mathbf{u}),\,j<i}\!\big(1-\hat{\alpha}_{j,t}(\mathbf{u})\big)\Big)\,\hat{\alpha}_{i,t}(\mathbf{u})\,\mathbf{c}_{i},$
where $\mathcal{S}_{t}(\mathbf{u})$ denotes the depth-sorted splats overlapping pixel $\mathbf{u}$. Parameters $\Theta=\{\bm{\mu}_{i},\bm{\Sigma}_{i},\alpha_{i},\mathbf{c}_{i}\}$ are optimized with the standard 3DGS photometric loss across frames.
Human Gaussians. We learn deformable human Gaussian representations from multi-view images that can be animated with different SMPL poses. Our approach consists of three key steps:
Step 1: Canonical Gaussian parameterization on SMPL. Given multi-view images of a person performing diverse poses, we learn a mapping from SMPL poses to 3D Gaussians in posed space. Following (li2023animatable), we place Gaussians on the 2D manifold of the SMPL surface by constructing an approximate front–back UV atlas via orthographic projections of the SMPL mesh. Let $\bm{\beta}$ denote SMPL shape and $\bm{\theta}_{t}$ the pose at time $t$. We rasterize pose-conditioned features into a pose map $P_{t}\in\mathbb{R}^{H\times W\times C}$ - denoted using $\mathcal{M}(\bm{\beta},\bm{\theta_{t}})$. A per-identity StyleUNet $f_{\phi}$ predicts a set of canonical human Gaussians anchored on the SMPL surface: $\mathcal{G}^{\mathsf{C}}_{t}\;=\;f_{\phi}(P_{t})\;=\;\big\{(\mathbf{x}_{k}^{\mathsf{C}},\bm{\Sigma}_{k}^{\mathsf{C}},\mathbf{c}_{k},\alpha_{k})\big\}_{k=1}^{N_{H}}.$
Step 2: Skinning to posed space (LBS). We obtain posed Gaussians by applying linear blend skinning (LBS) to canonical Gaussians $\mathcal{G}^{C}_{t}$ with SMPL joint transformations $\{(\mathbf{R}_{b}(\bm{\theta}_{t}),\mathbf{t}_{b}(\bm{\theta}_{t}))\}_{b=1}^{B}$ and vertex/bone weights $w_{kb}$ inherited from the SMPL surface by using nearest neighbour sampling from Canonical Gaussians to SMPL vertices. With $\mathcal{G}^{\mathsf{P}}_{t}=\{(\mathbf{x}_{k}^{\mathsf{P}},\bm{\Sigma}_{k}^{\mathsf{P}},\alpha_{k},\mathbf{c}_{k})\}_{k=1}^{N_{H}}$ we denote the posed human Gaussians at time $t$. During training, we render these posed Gaussians using standard 3DGS and supervise using multi-view images and cameras.
Step 3: Pose-to-Gaussian inference (test-time). Following (li2023animatable), we compute the top $K\!\in\![10,20]$ PCA components of training pose maps $\{P_{t}\}$, yielding mean $\bar{P}$ and basis $Q_{K}$. At inference, poses synthesized by our Gaussian-aligned motion module (Sec. 3.2) are first projected to this subspace and then mapped to posed Gaussians: $\tilde{P}_{y}\;=\;\bar{P}+Q_{K}z_{y}\qquad\mathcal{G}^{\mathsf{P}}_{y}\;=\;\mathrm{LBS}_{\bm{\theta}_{y}}\!\big(f_{\phi}(\pi\!\big(\mathcal{M}(\beta,\bm{\theta}_{y})\big))\big).$ Here we use $\pi$ to denote the projection to the subspace operation and $y$ to indicate a test-time pose. For further details please see supplementary materials.
3.2 Gaussian-Aligned Motion Synthesis
We introduce a Gaussian-aligned motion module that synthesizes controllable human motion directly in 3DGS scenes. Our key novelty is twofold: (i) we deploy reinforcement learning (RL) in Gaussian space by deriving reliable scene cues from opacity-weighted projections (no meshes or paired human–scene data required); and (ii) we couple RL locomotion with a deterministic latent optimizer for precise, contact-sensitive transitions in 3DGS scenes.
Design overview. We reuse a strong latent motion backbone trained on large scale mocap dataset and add two 3DGS-specific controllers: an RL locomotion policy that navigates between waypoints while avoiding scene Gaussians, and a deterministic latent-space optimizer that executes short, fine-grained actions near targets (e.g., stop, sit, grasp) before returning control to RL. While this explicit decomposition is not typical in existing motion synthesis frameworks, we find it especially effective in 3DGS settings, as this allows us to exploit the fact that in 3DGS scenes much of the raw scene detail can be abstracted to (i) a set of paths for navigation and (ii) action points (e.g., sitting locations, grasping targets provided by an animator) at which specific behaviors are executed - thus allowing for scene-aware motion synthesis without human-scene paired data. Both submodules operate consistently in the latent space of a learned motion model (Zhao:DartControl:2025).
Latent motion backbone. We adopt a latent motion prior, following prior work (Zhao:DartControl:2025) trained on AMASS (BABEL:CVPR:2021; mahmood19amass). Specifically, the model learns a compact motion-primitive space with a transformer VAE trained on mocap data, and places a diffusion prior in this latent space. Given motion history $\mathbf{H}$ and a future motion segment $\mathbf{X}$, the encoder $\mathcal{E}$ outputs a Gaussian posterior $q_{\phi}(\mathbf{z}\mid\mathbf{H},\mathbf{X})=\mathcal{N}(\bm{\mu},\sigma^{2}\mathbf{I})$ with reparameterized sample $\mathbf{z}=\bm{\mu}+\sigma\odot\bm{\varepsilon}$ where $\bm{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The decoder $\mathcal{D}$ reconstructs motion as $\widehat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\mathbf{z})$. On this latent space, a denoiser $\mathcal{G}$ operates with forward process
$q(\mathbf{z}_{\tau}\mid\mathbf{z}_{\tau-1})=\mathcal{N}(\sqrt{1-\beta_{\tau}}\,\mathbf{z}_{\tau-1},\,\beta_{\tau}\mathbf{I})$
and predicts the clean code $\widehat{\mathbf{z}}_{0}=\mathcal{G}(\mathbf{z}_{\tau},\tau,\mathbf{H},\mathbf{c})$, where $\mathbf{c}$ is an optional text embedding. During inference we sample $\mathbf{z}_{\tau_{\max}}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$, perform about $\tau_{\max}\!\approx\!10$ denoising steps to obtain $\widehat{\mathbf{z}}_{0}$, decode $\widehat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\widehat{\mathbf{z}}_{0})$, and update $\mathbf{H}$ with the last $H$ frames for autoregressive sampling. This latent backbone is reused; our contribution lies in coupling it with 3DGS-specific controllers.
Scene-adapted RL locomotion in 3DGS. Our insight here is that locomotion policies trained in mesh-based synthetic environments (zhao2023dimos) can be used in 3DGS reconstructions when combined with our scene adaptation. We cast navigation as an MDP whose action space is the latent space of the motion model. The policy outputs a start-noise $\mathbf{z}^{(\tau_{\max})}_{\text{RL},i}$, which a frozen $\mathcal{G},\mathcal{D}$ map to a short motion clip, ensuring stable rollouts. At step $i$, the agent observes state $s_{i}=(\mathbf{H}_{i},\mathbf{g}_{i},\bm{\eta}_{i},\mathbf{c}_{i})$ where $\mathbf{H}_{i}$ is motion history, $\mathbf{g}_{i}$ a goal cue, $\bm{\eta}_{i}$ a scene cue, and $\mathbf{c}_{i}$ a text embedding. The policy samples $a_{i}\sim\pi_{\theta}(\cdot\mid s_{i})$, interpreted as $\mathbf{z}^{(\tau_{\max})}_{\text{RL},i}$. The resulting clip $\mathbf{X}_{i}$ updates the history $\mathbf{H}_{i+1}$. Rewards $r_{i}=r(s_{i},a_{i},s_{i+1})$ encourage waypoint progress, obstacle avoidance, and kinematic plausibility. Training follows synthetic mesh-based environments as in (zhao2023dimos), while our contribution is the deployment in 3DGS. For deployment in 3DGS scenes (which lack meshes), we approximate navigation regions via orthographic projection: (i) compute PCA of Gaussian centers to align a top-down view, (ii) threshold opacities to filter floaters, (iii) render a binary map of obstacles and run A* for pathfinding. The policy consumes an egocentric occupancy grid/walkability map $\mathcal{M}\in\{0,1\}^{N\times N}$ centered on the agent. For each grid cell $u$, we compute its nearest-neighbor distance $d(u)$ to filtered Gaussians and mark $\mathcal{M}(u)=1$ if $d(u)>\tau$, else 0. Despite being approximate, this provides sufficiently reliable local context for navigation in 3DGS scenes. For inference during locomotion, we fix text cue $\mathbf{c}_{i}$ to ”walk”. For details please see supp mat.
Latent optimization for transitions in 3DGS. Once the agent reaches the vicinity of an action point, control switches from RL to deterministic latent-space optimization for fine-grained actions such as stopping, sitting, or grasping. Following (Zhao:DartControl:2025), we adopt a deterministic DDIM sampler (no step-skipping), which defines a fixed rollout (see supp. mat.) $\mathbf{M}=\mathrm{ROLLOUT}(\mathbf{Z}_{\text{opt}},\mathbf{H}_{\mathrm{seed}},\mathbf{C})$, where $\mathbf{Z}_{\text{opt}}$ is the terminal noise variable, $\mathbf{H}_{\mathrm{seed}}$ the seed history, and $\mathbf{C}$ a fixed text cue (“sit”, “grab”). We optimize $\mathbf{Z}_{\text{opt}}$ by minimizing
$\mathcal{L}(\mathbf{Z}_{\text{opt}})=F(\Pi(\mathbf{M}),\mathbf{g_{user}})+\mathrm{Cons}(\mathbf{M})$
with gradient updates $\mathbf{Z}_{\text{opt}}^{(k+1)}=\mathbf{Z}_{\text{opt}}^{(k)}-\eta\nabla_{\mathbf{Z}_{\text{opt}}}\mathcal{L}(\mathbf{Z}_{\text{opt}}^{(k)})$, where $\Pi(\cdot)$ projects the rollout onto task-relevant variables, $F$ measures goal satisfaction, and $\mathrm{Cons}$ adds continuity, collision, and smoothness constraints.
For position-only goals $\mathbf{g_{user}}$ (e.g., sitting or grabbing at a user-provided point), we synthesize a short $f$-frame snippet $\mathbf{M}=(\mathbf{M}_{1},\dots,\mathbf{M}_{f})$ starting from the locomotion end state $\mathbf{M}^{\mathrm{loc}}_{\mathrm{end}}$. In this setting, $F$ corresponds to the reach and stop terms, with $\mathcal{L}_{\mathrm{reach}}=\|\mathbf{x}_{f}(j^{\star})-\mathbf{g}\|_{2}^{2}$ and $\mathcal{L}_{\mathrm{stop}}=\|\mathbf{v}_{f}(j^{\star})\|_{2}^{2}$, while $\mathrm{Cons}$ corresponds to start-continuity $\mathcal{L}_{\mathrm{start}}=\|\mathbf{M}_{1}-\mathbf{M}^{\mathrm{loc}}_{\mathrm{end}}\|_{2}^{2}$, collision $\mathcal{L}_{\mathrm{coll}}=\sum_{b\in\mathcal{B}_{f}}[-\phi(b)]_{+}^{2}$, and smoothness $\mathcal{L}_{\mathrm{smooth}}=\tfrac{1}{f-1}\sum_{t=2}^{f}\|\mathbf{M}_{t}-\mathbf{M}_{t-1}\|_{2}^{2}$. The final objective is therefore
$\mathcal{L}=\mathcal{L}_{\mathrm{reach}}+\lambda_{v}\mathcal{L}_{\mathrm{stop}}+\lambda_{\mathrm{start}}\mathcal{L}_{\mathrm{start}}+\lambda_{\mathrm{coll}}\mathcal{L}_{\mathrm{coll}}+\lambda_{s}\mathcal{L}_{\mathrm{smooth}}$
Here $j^{\star}$ is an anchor joint (e.g., pelvis), $\mathbf{x}_{f},\mathbf{v}_{f}$ are its pose and velocity at frame $f$, $\mathcal{B}_{f}$ are sampled SMPL points at frame $f$, and $\phi(\cdot)$ is a differentiable signed-distance proxy to 3DGS Gaussians . After completing the action, the same formulation synthesizes the exit transition (e.g., sit $\to$ walk), after which locomotion resumes. Though our method can synthesize actions which mimic “picking up”, or “grabbing” in the scene it cannot lift actual objects from the scene - as we assume that the scene remains static throughout.
Figure 2: With (left) and right (without) refinment of Gaussians
3.3 Differentiable Contact Refinement in 3DGS
After animating the reconstructed human Gaussians with synthesized human motion data (Sec. 3.2), we place them into the reconstructed 3DGS scene. Here we want to highlight that we never use the SMPL mesh. We use the “SMPL pose” (the joint angles of the 24 SMPL joints or the 54 SMPL-X joints) - to drive the 3DGS avatar. A naive composition of posed human Gaussians with scene Gaussians often leads to floor/geometry penetration and inconsistent contacts. We introduce a contact-aware refinement that solves for small, physically meaningful translations of a sparse set of human Gaussians so that contacts are respected and penetrations are reduced. (Fig. 2)
We first note that mapping SMPL pose to 3D Gaussians is fundamentally different from mapping SMPL pose to 3D SMPL mesh. Mapping SMPL pose to a mesh is just a linear blend skinning operation that maps canonical SMPL mesh vertices to posed SMPL vertices - hence the SMPL pose can usually be easily optimized to maintain contact with the scene. This is not the case with 3D human Gaussians. Mapping SMPL pose to 3D human Gaussians involves a forward pass through the learnt network - and then applying LBS to the output Gaussians. If we were to optimize the SMPL pose for correct contact this would entail getting a gradient through the neural network that maps SMPL pose to 3D Gaussians - which would probably be difficult to converge. Instead we optimize per-Gaussian offsets as the Gaussians (after being output by the network) are already placed close enough to reasonable locations in the 3d scene; thus we can simply optimize per gaussian offsets - with heavy regularization for correct contact. We describe the setup in detail below.
Setup. Let the posed human Gaussians at time $t$ be $\mathcal{G}^{\mathsf{P}}_{t}=\{(\mathbf{x}_{k}^{\mathsf{P}},\bm{\Sigma}_{k}^{\mathsf{P}},\alpha_{k},\mathbf{c}_{k})\}_{k=1}^{N_{H}}$, and the scene Gaussians be $\mathcal{G}^{S}$. Our goal is to refine a subset of the human Gaussians by per-frame translations $\mathbf{T}_{k,t}$ to achieve (i) contact where appropriate and (ii) separation elsewhere.
Contact detection and indexing. From synthesized SMPL motion, we detect contact frames for a set of body joints using simple kinematic cues. For joint $c$ with position $\mathbf{p}_{c,t}$, velocity $\mathbf{v}_{c,t}=\mathbf{p}_{c,t}-\mathbf{p}_{c,t-1}$ and acceleration $\mathbf{a}_{c,t}=\mathbf{v}_{c,t}-\mathbf{v}_{c,t-1}$, a frame is marked as contact if $\delta_{c,t}=(|v_{c,t}^{y}|<\tau_{v})\wedge(a_{c,t}^{y}<\tau_{a})$, where $y$ is the vertical axis. Because human Gaussian templates have identity-dependent counts and no global correspondence, we lift SMPL contact vertices $V_{c}^{\text{SMPL}}$ (e.g., feet, hip) to the human Gaussians via nearest-neighbour search in the canonical space: $i^{\star}=\arg\min_{k}\|\mathbf{x}_{k}^{\mathsf{C}}-\mathbf{u}\|_{2}$, $\mathbf{u}\in V_{c}^{\text{SMPL}}$. The resulting index set $\mathcal{I}_{c}$ specifies which human Gaussians may be refined at contact.
Scene proximity in Gaussian space. We measure scene proximity using a soft nearest-neighbour distance to scene Gaussians
$d_{\beta}(\mathbf{x})=-\tfrac{1}{\beta}\log\!\Big(\sum_{j=1}^{N_{S}}\exp\big(-\beta\,\|\mathbf{x}-\bm{\mu}_{j}\|\big)\Big),$
where $\bm{\mu}_{j}$ are scene Gaussian centers and $\beta$ controls softness. This provides stable gradients for contact/separation without requiring explicit meshes.
Refinement objective. For a contact Gaussian $k\in\mathcal{I}_{c}$ at frame $t$ with indicator $\delta_{c,t}$, we optimize a translation $\mathbf{T}_{k,t}$ and update $\tilde{\mathbf{x}}_{k,t}^{\mathsf{P}}=\mathbf{x}_{k,t}^{\mathsf{P}}+\mathbf{T}_{k,t}$ by minimizing
$\mathbf{T}_{k,t}^{\star}=\arg\min_{\mathbf{T}}\;\lambda_{s}\,\|\mathbf{x}_{k,t}^{\mathsf{P}}+\mathbf{T}-\bm{\mu}_{j(k,t)}\|_{2}^{2}\;+\;\lambda_{d}\,\psi(\mathbf{x}_{k,t}^{\mathsf{P}}+\mathbf{T},\delta_{c,t})\;+\;\lambda_{r}\,\|\mathbf{T}\|_{2}^{2},$
where $\bm{\mu}_{j(k,t)}$ is the nearest scene Gaussian center and
$\psi(\mathbf{x},\delta)=\begin{cases}d_{\beta}(\mathbf{x})^{2},&\delta=1\;\text{(enforce contact)}\\ h_{r}\big(d_{\beta}(\mathbf{x})\big)^{2},&\delta=0\;\text{(enforce separation)}\end{cases}\quad\text{with}\;h_{r}(d)=\max(0,\,r-d).$
For temporal coherence, we add $\lambda_{t}\sum_{t}\|\mathbf{T}_{k,t}-\mathbf{T}_{k,t-1}\|_{2}^{2}$. Intuitively, the objective snaps designated contact Gaussians toward nearby scene surfaces when contact is detected, pushes them away otherwise, penalizes large displacements, and smooths motion over time.
The refined human Gaussians $\tilde{\mathcal{G}}^{\mathsf{P}}_{t}=\{(\tilde{\mathbf{x}}_{k,t}^{\mathsf{P}},\bm{\Sigma}_{k}^{\mathsf{P}},\alpha_{k},\mathbf{c}_{k})\}_{k=1}^{N_{H}}$ are composed with $\mathcal{G}^{S}$ and rendered with the standard 3DGS rasterizer to produce photorealistic interactions (e.g., walking, sitting) with improved contact fidelity and fewer penetrations. To the best of our knowledge, this is the first mesh-free refinement in Gaussian space that leverages a differentiable scene-distance, remains identity-agnostic via SMPL-to-Gaussian lifting, and operates as a lightweight post-hoc stage to improve contact realism without retraining.
Figure 3: Qualitative results: Our method generates diverse motions across scenes and identities.
4 Experiments
For rendering evaluation, we present two modified mesh-based baselines (Baseline A and B) and evaluate with two evaluation protocols (I-human and II-automated). For further evaluation on motion quality and ablations please see the supplementary.
Table 1: Evaluation design. Two baselines $\times$ two protocols. HQ: highest-quality rendering settings for each method. The same camera trajectories are used within each pairwise comparison.
Setting Dataset / Source 3DGS Scene (ours) Recons Mesh Scene (Baseline) Baseline Rendering Protocols
Baseline A Mon. Vids (same scenes) 3DGS reconstruction VGGT dense 3DGS I and II
Baseline B Replica and Curated 3DGS SuperSplat Replica Mesh I and II
4.1 Rendering Evaluation
Baseline A: For Baseline A we collect monocular videos from DL3DV (ling2024dl3dv); each scene is reconstructed twice (once as 3DGS, once as a mesh using dense VGGT reconstruction (wang2025vggt)) so that comparisons are within-scene. Using the meshes obtained using dense VGGT reconstruction, we again use (zhao2023dimos) to generate SMPL-X parameters. Then we use these parameters naively to pose human Gaussians (Sec. 3.1) in the 3D scene and render the composited scene and human Gaussians using 3DGS. Note we do not perform any refinement. Furthermore note that the scene mesh is only used for motion synthesis but for rendering we use the 3DGS scene reconstruction and the posed human Gaussians. Baseline A is designed to show that a naive baseline that composites human and scenes Gaussians does not work out-of-the-box for monocular videos and hence provides further motivation for our algorithm. For baseline A evaluations, we render synchronized camera trajectories per pair (identical poses, FoV, and exposure).
Baseline B: In this experiment we aim to evaluate the rendering quality of a strong mesh-based baseline vs our 3DGS based algorithm. We use the highest quality existing mesh based 3D scenes from the Replica (replica19arxiv) dataset. In the Replica Scene we use the framework in (zhao2023dimos) to synthesize motion. Then we use a rigged scan from RenderPeople (along with its texture map) (RenderPeople) animated with the synthesized motion parameters in the Replica scene to generate the final renderings. For rendering our videos we use scenes from the Supersplat library and Avatars from Avatarrex (zheng2023avatarrex) dataset. This experiment aims to evaluate the highest quality rendering of a mesh based rendering vs a highest quality 3DGS renderings for the specific setting of human-scene animation. Note comparisons are not within scene.
Here we want to highlight that finding the same scene+human combination for both our method and Baseline B would be difficult - because the data capture pipelines for 1) 3DGS vs mesh scene 2) 3DGS vs mesh human are fundamentally different. RenderPeople uses a rig of 250 DSLR cameras to capture a 3D human mesh - while the AvatarX dataset only uses 16 cameras and Behave dataset uses only 4. For 3DGS Avatar reconstruction motion of about 120 seconds inside the multiview camera setup is required while for a mesh capture only one frame is required. The way 3D scenes are reconstructed for Replica (mesh) and SuperSplat (3DGS) are also fundamentally different - the Replica reconstruction uses depth while some Supersplat scenes use lidar. Additionally the Replica original images are not available so we can’t reconstruct a 3DGS splat for Replica scenes of the same quality as the ones on SuperSplat. As such for this particular baselines comparisons are not within scene.
We also want to acknowledge that for 3DGS animatable Avatars we require multiview video while a mesh can be reconstructed using multiview images. However we believe that for different use-cases, users would be willing to make the tradeoff for higher rendering quality.
Evaluation Protocol I: Human preference study We conduct a pairwise forced-choice study measuring perceived photorealism. Each trial presents two mute videos from ours vs Baseline A or ours vs Baseline B. Participants select the video they find more photorealistic. We generate 5 samples for both comparisons and aggregate votes by comparison. We collect 21 participants for both Baseline A and B. We report win rate (%) of Ours over its comparator.
Table 2: Human preference study (win rate, %) — fraction of pairwise trials where Ours is preferred. Baseline B compares Ours vs a mesh based baseline at highest-quality; Baseline A compares Ours (3DGS) vs a custom baseline designed for monocular videos. Higher is better.
Replica vs 3DGS-Library (Baseline B) Monocular (Baseline A)
Ours(3DGS)  vs Mesh 82.1 72.9
Evaluation Protocol II: VLM-based pairwise judgment We use two strong vision–language model (VLM): GPT-5 (openai2025gpt5) and Gemini2.5 (comanici2025gemini25pushingfrontier) as paired comparators between still renderings. For each video pair, we uniformly sample 10 frames per method, form matched pairs at the same timestamps, and query the VLM with “which image looks more photoreal?”. The VLM outputs a ternary judgment {Left wins, Right wins, Tie}; we compute a VLM win rate as the percentage of non-tied pairs favoring Ours. We randomize image order and prevent leakage by removing textual overlays and metadata.
For the pairwise VLM study, the VLM is instructed to ignore artistic style and focus on physical plausibility: “which image looks more photoreal? Consider geometry (straight lines, depth cues), materials (BRDF, speculars), lighting/shadows, and absence of artifacts (flicker, halos, floaters).
Table 3: VLM preference study (win rate, %) — fraction of pairwise comparisons where Ours is preferred. Baseline B compares Ours vs a mesh based baseline at highest-quality; Baseline A compares Ours (3DGS) vs a custom baseline designed for monocular videos. Higher is better.
Replica vs 3DGS-Library (Baseline B) Monocular (Baseline A)
GPT-5 75.2 71.8
Gemini 2.5 69.1 65.9
Results Baseline A (Same monocular video) On the within-scene comparison (Fig. 5), Ours outperforms the mesh baseline in both human and VLM judgments (Tables 23). Note that the scene meshes reconstructed using VGGT often exhibit blocky strucutres, blocked paths and hence are not suitable for motion synthesis, while our algorithm directly operates in Gaussian Space and doesnt suffer from the same problems. For detailed failure modes see Supp Mat.
Baseline B (Replica Scene), HQ vs HQ. Across both evaluations (Fig. 5) Ours is preferred by humans and by the VLM comparator (Tables 23) - thus clearly underscoring the central premise of the paper - that neural scene representations yield better rendering quality for human-scene interaction compared to existing mesh based representations.
Figure 4: Free viewpoint rendering of edited monocular video with animated humans
[…]
Figure 5: Visual Comparisons with Baselines
4.2 Qualitative Results and Free viewpoint rendering of Edited Videos
In Fig. 3, we show results for diverse scenes from the SuperSplat library, Scannet scenes with Avatars from BEHAVE (sparse only 4 cameras), DNA-Rendering (cheng2023dnarendering), Avatarrex (zheng2023avatarrex) datasets. In Fig. 4, we demonstrate that our method works for monocular RGB scene videos and allows for free viewpoint rendering of videos edited with geometry consistent placement of animated humans in the scene. For more results please see the supplementary.
5 Conclusion
We have presented, to the best of our knowledge, the first method to synthesize human interactions in diverse 3D environments using 3D Gaussian Splatting (3DGS) as the underlying 3D representation. Our results suggest that neural rendering is now mature enough to function as a practical component in end-to-end 3D human–scene animation pipelines, bridging previously disjoint lines of work in human-scene animation and neural rendering. Crucially, our pipeline operates on scenes reconstructed from monocular RGB video and allows for applications such as monocular RGB geometry consistent video editing. We believe this framing and evidence open new research directions at the intersection of human animation, scene understanding, and neural rendering.
Limitations. Despite this progress, our pipeline has several limitations. First, complex and rapidly changing illumination can cause rendering artifacts and imperfect relighting. Second, we do not enforce full physics-based constraints, which can yield interactions that look plausible yet violate contact, stability, or momentum conservation. Third, the range of interaction types is limited; highly dexterous manipulation and long-horizon, multi-contact behaviors remain challenging. Fourth, we assume access to multiview videos of a human performing diverse actions.
Outlook. Addressing these issues suggests several promising research directions: integrating stronger lighting estimation and inverse rendering, incorporating differentiable or learned physics priors for contact and dynamics, expanding the interaction vocabulary to richer, longer, and multi-person scenarios and investigation how Avatars that generalize to Out-of-distribution poses can be reconstructed from monocular videos. We hope this work provides a foundation for scalable, video-native human–scene animation pipelines and catalyzes further advances in data, models, and evaluation for interactive 3D human animation.